# FENDA-FL: Personalized Federated Learning on Heterogeneous Clinical Datasets

### v2

@@ Abstract
Federated learning (FL) is increasingly being recognized as a key approach to overcoming the data silos that so frequently obstruct the training and deployment of machine-learning models in clinical settings. This work contributes to a growing body of FL research specifically focused on clinical applications along three important directions. First, we expand the FLamby benchmark (du Terrail et al., 2022a) to include evaluation of personalized FL methods and demonstrate substantive performance improvements over the original results. Next, we advocate for a comprehensive checkpointing and evaluation framework for FL to reflect practical settings and provide multiple comparison baselines. Finally, we study an important ablation of PerFCL (Zhang et al., 2022). This ablation is a natural extension of FENDA (Kim et al., 2016) to the FL setting. Experiments conducted on the FLamby benchmarks and GEMINI datasets (Verma et al., 2017) show that the approach is robust to heterogeneous clinical data and often outperforms existing global and personalized FL techniques, including PerFCL.
Data and Code Availability
In this work, two real-world clinical dataset collections are used. Three datasets are drawn from the public FLamby benchmarks (du Terrail et al., 2022a). [^1] Code and documentation for FLamby is found at https://github.com/owkin/FLamby All code associated with the FLamby experiments, including the scripts to perform hyper-parameter sweeps, selection, and evaluation is publicly released. [^2] All experimental code is available at: https://github.com/vectorInstitute/FL4Health Artifacts generated by the FLamby experiments, including checkpoints, are available upon request. Two datasets are sourced from the GEMINI consortium, which are derived from up-to-date patient data at Canadian hospitals (Verma et al., 2017). Due to the sensitive nature of the data and GEMINI’s closed environment, the datasets and code are not publicly released. However, we are happy to share code with individuals who have access to GEMINI. [^3] GEMINI data access may be requested at https://www.geminimedicine.ca/access-data
Institutional Review Board (IRB)
IRB information will be provided upon publication.
@@ 1 Introduction
Federated learning (FL), first introduced by McMahan et al. (2017), provides an avenue for training models in distributed data settings without requiring training data transfer. The focus of this work is cross-silo FL (Kairouz et al., 2021), where participants represent a small number of reliable institutions with sufficient computing resources. In particular, we focus on the challenging setting of distributed heterogeneous datasets. Such datasets occur in many FL settings and are particularly common in clinical environments, as the statistical properties of local datasets are influenced by, for example, disparate patient populations and varying physical equipment.
Heterogeneous, or non-identically and independently distributed (non-IID), datasets often require special treatment in FL frameworks to avoid issues with convergence and to improve performance. This work provides an extensive evaluation of state-of-the-art personalized and non-personalized FL methods in the context of real-world clinical data. As part of the experimental design, we advocate for a more comprehensive evaluation framework for FL methods. This includes utilizing reasonable checkpointing strategies, bringing FL training into better alignment with standard ML practices and evaluating model performance from several perspectives. Such a framework aims to improve the evaluation of new FL techniques and better reflect the settings of practical deployment.
Two clinically relevant dataset collections are considered, the public FLamby benchmark (du Terrail et al., 2022a) and GEMINI (Verma et al., 2017), a large Canadian clinical data consortium. The experiments represent a significant extension of the FLamby benchmark along two dimensions. First, the results establish marked improvements in the performance of many FL approaches for the three tasks considered. Second, this work expands the benchmark results to include an evaluation of personal FL models, which often further improve performance.
Finally, a novel ablation of the recently proposed PerFCL method (Zhang et al., 2022) is considered. The ablation results in a natural extension of the frustratingly easy neural domain adaptation (FENDA) method (Kim et al., 2016) to the FL regime and, thus, is termed FENDA-FL. The approach falls under the category of Personalized FL (Tan et al., 2022), where each client trains a model uniquely suited to its local distribution. The results to follow demonstrate that FENDA-FL often outperforms existing FL techniques, including PerFCL, despite being simpler. Moreover, it has the advantage of a more flexible architecture, admitting inductive bias through intentional model design to further improve performance.
2 Related Work and Contributions
FL has been successfully applied in many contexts (Hard et al., 2018; Wang et al., 2019; Dupuy et al., 2022). Moreover, it remains an active area of research along many dimensions, from communication efficiency (Passban et al., 2022; Isik et al., 2023) to improved optimization (Reddi et al., 2021). While the number of FL applications in clinical settings is growing (du Terrail et al., 2022b; Dayan et al., 2021), the approach is still relatively under-utilized, despite its potential benefits. This is due, in part, to a dearth of research and tooling focused on this important area. This paper aims to deepen fundamental understanding of how FL may be used to overcome clinical data silos and train models with potentially greater performance than even centralized data settings.
The study of FL in clinical settings has been expanding. Some works focus on the utility of FL for specific clinical applications (Andreux et al., 2020b, a; Chakravarty et al., 2021; Wang et al., 2023; Gunesli et al., 2021). The FLamby benchmark (du Terrail et al., 2022a) aims to standardize the datasets in these studies to facilitate systematic evaluation of FL in healthcare. We leverage several datasets from the FLamby collection, improving upon many of the originally reported baselines and establishing new performance benchmarks for personalized FL methods. We also extend the understanding of the practical use of FL in clinical settings by measuring results on datasets from the GEMINI consortium. Finally, we release an open-source library aimed at making FL research and evaluation easy and reproducible.
A set of work aimed at training personalized FL models already exists (Tan et al., 2022; Fallah et al., 2020; Yu et al., 2020; Liang et al., 2020). These approaches aim to federally train models that produce elevated performance on each participating clients’ individual data distribution. Several state-of-the-art personalized FL techniques, FedPer (Arivazhagan et al., 2019), APFL (Deng et al., 2020), and PerFCL (Zhang et al., 2022), are extensively evaluated herein. FENDA-FL performs comparably or better in most settings. Such experiments extend the FLamby benchmark to include the evaluation of personalized FL approaches, which generally improve upon non-personal FL techniques, including those aimed at heterogeneous data challenges. Some methods, (Li et al., 2021b; Jeong and Hwang, 2022; Peterson et al., 2019), consider narrow settings or architectures and are not considered.
FENDA-FL is connected to Domain Adaptation (DA). Generally, DA involves learning a model from source datasets that generalizes to target datasets. During training, subsets of the target datasets, typically unlabeled or sparely labeled, are available. Federated DA considers the setting of distributed source and target datasets. A server may host the source dataset coupled with client-based, unlabelled datasets (Yao et al., 2022) or vice-versa (Peng et al., 2019). Alternatively, both source and target datasets may be distributed (Shen et al., 2023; Song et al., 2020). Relatedly, Federated Domain Generalization focuses on the task of training a model on distributed, private source domains that generalizes to unseen target domains (Li et al., 2023). Such methods focus on training models for out-of-distribution clients rather than (personalized) models optimized for known client data. The settings above are related to those considered here but have different objectives and data settings. However, FENDA-FL establishes a novel conceptual bridge between FL and DA.
Finally, an important gap in the current literature surrounding FL is a set of guidelines for evaluating FL models that closely resemble practical deployment. Many works (Karimireddy et al., 2020; McMahan et al., 2017; Li et al., 2021a; du Terrail et al., 2022a) simply perform a fixed number of server rounds and compare the resulting models along a small set of dimensions. This work of details effective checkpointing strategies and a comprehensive set of criteria required to demonstrate robust model performance compared with baseline approaches that do not leverage federated training.
@@ 3 Methodology
As in (du Terrail et al., 2022a), the experiments to follow incorporate a number of existing FL techniques as baselines along with a number of more recent approaches, including FENDA-FL. The methods evaluated are FedAvg (McMahan et al., 2017), FedAdam (Reddi et al., 2021), FedProx (Li et al., 2020), SCAFFOLD (Karimireddy et al., 2020), MOON (Li et al., 2021a), FedPer (Arivazhagan et al., 2019), APFL (Deng et al., 2020), and PerFCL (Zhang et al., 2022). Two non-FL training setups are also considered. The first, central training, pools all data to train a single model. The second is local training, where each client trains a model solely on local data, resulting in a model for each client with no global information.
@@
The original FedAvg algorithm, which applies weighted parameter averaging, remains a strong FL baseline. FedAdam is a recent extension of FedAvg that incorporates server-side first- and second-order momentum estimates and has shown some robustness to data drift. In all experiments, $\beta_{1}=0.9$, $\beta_{2}=0.99$, and $\tau=1\mathrm{e}{\text{-}9}$. The FedProx and SCAFFOLD methods aim to address issues associated with non-IID datasets by correcting local-weight drift through a regularization term or control variates, respectively. MOON builds on these ideas, applying a contrastive loss to constrain drift in the client models’ feature representations at a selected layer. Each method trains a global set of weights, updated at each server round, shared by all clients.
@@
For the personalized FL baselines, each client possesses a unique model. FedPer splits the models into a feature extractor and classifier. The feature extractor is shared among clients and federally trained, while the classifier is unique to each client. In the APFL approach, predictions are made through a convex combination of twin models. One is federally trained using FedAvg; the other incorporates only local updates. The combination parameter, $\alpha$, is adapted during training using the gradients of the global and local models. Finally, PerFCL extends MOON to consider two feature extractors, one is shared by all clients and the other is specific to each client along with a local classification layer. The feature extractors are constrained using different contrastive losses.
For all experiments, clients use an AdamW optimizer (Loshchilov and Hutter, 2018) with default parameters for local training. The only exception is SCAFFOLD, which explicitly assumes an unmodified learning rate is applied and, thus, uses standard SGD. Hyper-parameters for each method, summarized in Table LABEL:hyper-parameter_summary, are tuned. The ranges tested and best hyper-parameters for each method and dataset are detailed in Appendices A and B, along with the selection methodology.
@@
hyper-parameter_summary Method Param 1 Param 2 Param 3 FedAvg Client LR FedAdam Client LR Server LR FedProx Client LR $\mu$ SCAFFOLD Client LR Server LR MOON Client LR $\mu$ FedPer Client LR APFL Client LR $\alpha$ LR PerFCL Client LR $\mu$ $\gamma$ FENDA-FL Client LR
Table 1: A summary of hyper-parameters tuned for each method. LR denotes learning rate.
3.1 FENDA-FL
For PerFCL, the authors consider a model with parallel global and local feature extractors feeding into a locally trained classifier. The latent spaces of the feature extractors are constrained with contrastive loss functions. These loss functions are weighted with fixed hyper-parameters $\mu$ and $\gamma$. Experiments with the approach show that PerFCL improves upon MOON and other FL baselines. However, therein, PerFCL is not compared against other personalized methods. Moreover, no ablation studies are considered to quantify the benefit associated with the architecture compared with the contrastive loss. As such, we study the effect of removing the contrastive losses.
Removal of the contrastive losses results in an approach that directly extends the FENDA method (Kim et al., 2016), originally designed for domain adaptation, to FL. In FENDA, each domain associated with a dataset has two feature extractors. One is exclusively trained on domain-specific data, while the other is shared and trained across all domains. The extracted features are then combined and processed by a domain-specific classification head. The shared feature extractor is free to learn robust, domain-agnostic features, while the domain-specific extractor learns features important in making accurate local predictions. Should the global features not prove useful for domain-specific prediction, the classification head can simply ignore those features and make predictions based solely on local features.
As with PerFCL, each FENDA-FL client model consists of a global and local feature extraction module. The two sets of activations are concatenated and processed by a local classification network. Each client trains the full network with local data during the client training phase. Thereafter, the clients only transmit the weights of the global feature extractors for server-side aggregation, performed through FedAvg in the experiments below. An illustration of FENDA-FL is shown in Figure LABEL:fenda_fl_two_client and a technical discussion is found in Appendix J, which includes the forms of distribution drift handled by the method.
@@ 3.1 FENDA-FL
fenda_fl_two_client Refer to caption
Figure 1: An illustration of FENDA-FL. Each client’s model consists of local and global modules, and a local classifier. Global modules are exchanged with the server.
FENDA-FL has several advantages over PerFCL, beyond a simpler setup. Removal of the contrastive losses eliminates the need to tune two hyper-parameters. Further, it has been observed that similar loss induced constraints may hinder learning, especially when hyper-parameters are poorly calibrated. For example, Li et al. (2020) observe that the performance of FedProx is sensitive to the value of $\mu$. As seen in Appendix G, while APFL is somewhat robust to the choice of learning rate for $\alpha$, there is still a meaningful impact on performance. Without the contrastive losses, FENDA-FL is also a more flexible approach - global and local modules need not have the same latent space, allowing for the injection of inductive bias through architecture variation. As shown in Section 4 and Appendix E, this is useful in clinical settings to produce optimal performance.
3.2 Clinical Datasets
Five distinct clinical datasets are considered in the experiments. Three are drawn from the FLamby benchmark (du Terrail et al., 2022a) and two are privately held within the GEMINI consortium (Verma et al., 2017). Each represents a clinically relevant task and data derived from real patients. The selected FLamby datasets are Fed-Heart-Disease, Fed-IXI, and Fed-ISIC2019. Fed-Heart-Disease is a binary classification task predicting whether a patient has heart disease from a collection of measurements. Fed-IXI considers MR images of patients’ heads with the target of 3D segmentation of brain-tissue within the scans. Finally, Fed-ISIC2019 is a highly imbalanced, multi-class dermatology dataset consisting of 2D images. The task is to classify the images into one of eight different melanoma categories. More details on these datasets are found in Appendix A.
The two GEMINI datasets consist of clinical data from seven hospitals in Canada. The data is anonymized by removing patient identifiers, yet still contains highly sensitive patient-level information. The costs and risks associated with centralizing such data are among the main motivations for this work. Given that the GEMINI consortium covers a huge set of variables associated with patient encounters, feature selection and preprocessing are conducted to prepare the source data for each task. Additional details are found in Appendix B. The first task is in-hospital mortality prediction, which determines the probability of the patients’ latest encounter ending in mortality based on admission records. The second task aims to predict the risk of delirium in patients based on a wide variety of clinical measurements.
3.3 Federated Checkpointing and Evaluation
To address an important gap in the FL literature, we propose a straightforward checkpointing strategy and evaluation framework for FL models, building on the foundations in (du Terrail et al., 2022a) and better aligning with standard ML model practices. In many works (du Terrail et al., 2022a; Li et al., 2020; Karimireddy et al., 2020; Passban et al., 2022), federated training proceeds for a fixed number of server rounds. Thereafter, the final model is evaluated on held-out data, representing either a central or distributed test set. However, as with standard ML training, the model produced at the end of a training trajectory is not always optimal due to phenomena such as over-fitting.
We propose splitting each clients’ dataset into train and validation sets. In the experiments, the split used is 80–20. The validation splits enable two distinct checkpointing approaches. After each round of local training and aggregation, the clients evaluate the model on their local validation sets. In local checkpointing, each client uses their local validation loss to determine whether a new checkpoint should be stored. Alternatively, in global checkpointing, the client losses are aggregated by the server using a weighted average. This average is used to checkpoint a global model to be shared by all clients. Note that, in both cases, the best model is not necessarily the one produced by the final server round. In the local case, clients are free to keep the model they deem best. In the global strategy, all clients share the same model from some point in the server rounds. The personalized FL strategies do not have a single global model and only use local checkpointing.
To evaluate checkpointed models, we advocate for at least three separate baselines to be considered, two of which are part of the FLamby benchmarks. Models trained on central data are a common gold standard for FL performance. In the experiments, a single model is trained on data pooled from each of the clients and checkpointed based on a validation loss. The model is evaluated on each client’s test set and the performance uniformly averaged. In local training, a client trains a model on its own training data. The model is shared with all clients and evaluated on their test data, including that of the original client. This setting mimics the scenario where, for example, a hospital distributes a model to other hospitals. Again, the measured metrics are averaged across clients. An important alternative setting, referred to here as siloed, is one where each client trains its own model and evaluates on its local test data. The results are then averaged across the clients. This simulates a strong baseline where, for example, the problem is important enough that each hospital has collected its own data to train a model. If an FL model does not produce better performance, there is significantly less incentive to participate in collaborative training.
To quantify stability and smooth randomness in training, five FL training runs with distinct validation splits are performed for each method. Within each run, the performance across clients is averaged. The mean and $95$% confidence intervals for this average are reported. When local checkpointing is used, each client loads its respective model to be evaluated on that client’s test data.

### v3

# A Comprehensive View of Personalized Federated Learning on Heterogeneous Clinical Datasets

@@ Abstract
Federated learning (FL) is increasingly being recognized as a key approach to overcoming the data silos that so frequently obstruct the training and deployment of machine-learning models in clinical settings. This work contributes to a growing body of FL research specifically focused on clinical applications along three important directions. First, we expand the FLamby benchmark (du Terrail et al., 2022a) to include a comprehensive evaluation of personalized FL methods and demonstrate substantive performance improvements over the original results. Next, we advocate for a comprehensive checkpointing and evaluation framework for FL to reflect practical settings and provide multiple comparison baselines. To this end, an open-source library aimed at making FL experimentation simpler and more reproducible is released. Finally, we propose an important ablation of PerFCL (Zhang et al., 2022). This ablation results in a natural extension of FENDA (Kim et al., 2016) to the FL setting. Experiments conducted on the FLamby benchmark and GEMINI datasets (Verma et al., 2017) show that the proposed approach is robust to heterogeneous clinical data and often outperforms existing global and personalized FL techniques, including PerFCL.
@@ 1 Introduction
Federated learning (FL), first introduced by McMahan et al. (2017), provides an avenue for training models in distributed data settings without requiring training data transfer. The focus of this work is cross-silo FL (Kairouz et al., 2021), where participants represent a small number of reliable institutions with sufficient computing resources. In particular, we focus on the challenging setting of distributed heterogeneous datasets. Such datasets occur in many FL settings and are particularly common in clinical environments, as the statistical properties of local datasets are influenced by, for example, disparate patient populations and varying physical equipment. Such datasets often require special treatment in FL frameworks to avoid issues with convergence. Personalized FL methods have proven promising for such non-identically and independently distributed (non-IID) datasets.
Another meaningful challenge in the study of FL methods in clinical settings is the lack of access to high-quality benchmarking pipelines and real-world medical datasets. Therefore, we provide a thorough benchmark of state-of-the-art personalized and non-personalized FL methods, including a promising ablation, evaluated with a comprehensive framework on real-world public and private clinical datasets. The main contributions are as follows.
  • We advance existing work in personalized FL benchmarking to consider the challenging setting of heterogeneous clinical data. In so doing, marked improvements in the performance of many FL approaches are demonstrated for the datasets considered. This includes establishing that personalized methods offer significant advantages over non-personalized approaches for real-world clinical tasks.
  • We advocate for a more comprehensive evaluation framework for FL methods. This includes utilizing reasonable checkpointing strategies, bringing FL training into better alignment with standard ML practices, and evaluating model performance from several perspectives. With this work, an open-source library aimed at making FL research and evaluation easier and more reproducible is released. [^1] All experimental code is available at: https://github.com/vectorInstitute/FL4Health
  • A novel ablation of the recently introduced PerFCL algorithm (Zhang et al., 2022) is proposed that results in a natural extension of the frustratingly easy neural domain adaptation (FENDA) method (Kim et al., 2016) to the FL regime and, thus, is termed FENDA-FL. Such an ablation proves to be important as experiments demonstrate that the approach often outperforms existing FL techniques.
Generalizable Insights in the Context of Healthcare
The results in this work provide important insights and value to those interested in leveraging FL in clinical applications. The experiments provide a thorough evaluation of carefully selected personal and non-personal FL methods, as discussed in Appendix A, on real-world clinical tasks, demonstrating that personalized methods provide marked performance improvements. Notably, in most settings, personalized FL approaches even outperform models trained on centralized data. The evaluation framework put forward herein provides rigorous assessment of the studied FL methods, and the proposed checkpointing approaches ensure optimal performance. Finally, the proposed ablation of PerFCL, FENDA-FL, produces consistently strong results across most of the clinical tasks investigated, see Table 1, making it a compelling candidate for further application in healthcare.
2 Related Work
2.1 Standard and Personalized FL
FL has been successfully applied in many contexts (Hard et al., 2018; Wang et al., 2019; Dupuy et al., 2022) to jointly learn ML models on distributed data. It is an active area of research along many dimensions, from communication efficiency (Passban et al., 2022; Isik et al., 2023) to improved optimization (Reddi et al., 2021). A central challenge in FL is overcoming heterogeneous data distributions arising from variations in the underlying data generating process of clients. To address this, personalized FL approaches aim to federally train models that produce elevated performance on each participating client’s individual data distribution. Several state-of-the-art personalized FL techniques, FedPer (Arivazhagan et al., 2019), Ditto (Li et al., 2021b), APFL (Deng et al., 2020), and PerFCL (Zhang et al., 2022), are extensively evaluated herein. The FENDA-FL approach, discussed in detail below, performs comparably or better in most settings.
2.2 FL for Clinical Tasks and Benchmarking
While the number of FL applications in clinical settings is growing (du Terrail et al., 2022b; Dayan et al., 2021), the approach is still relatively under-utilized, despite its potential benefits. This is due, in part, to a dearth of research and tooling focused on this important area. Some existing works focus on the utility of FL for specific clinical applications (Andreux et al., 2020b, a; Chakravarty et al., 2021; Wang et al., 2023; Gunesli et al., 2021). The FLamby benchmark (du Terrail et al., 2022a) aims to standardize the datasets in these studies to facilitate systematic evaluation of FL in healthcare. While the benchmark has significant utility, the limited evaluation protocols and omission of personalized FL methods narrow the insights of the benchmark in the context of clinical applications. While there are some benchmarks focused on personalized FL methods (Matsuda et al., 2024; Chen et al., 2022), evaluation of personalized FL in the healthcare domain is unexplored.
Our work aims to provide a broader account of the performance of FL, including personalized methods, on clinical datasets using reasonable checkpointing and evaluation strategies that are more closely aligned with standard ML practices. To this end, we leverage several datasets from the FLamby collection, improving upon many of the originally reported baselines and establishing new performance standards. We also extend the understanding of the practical use of FL in clinical settings by measuring results on datasets from the GEMINI consortium. The experiments demonstrate that personalized FL methods generally improve upon non-personalized FL approaches and highlight the importance of intelligent checkpointing strategies.
@@ 3 Methodology
In this work, the performance and utility of non-personal and personal FL algorithms are measured. Non-personalized methods are important when a single model is desired across clients and can be effective, even in heterogeneous data settings. On the other hand, personalized FL techniques produce models that benefit from cross-client aggregation, while retaining client-specific representations. As such, these approaches provide a promising balance between global and local information and are of specific interest in clinical settings, where performance is of the utmost importance.
@@
As in du Terrail et al. (2022a), the experiments to follow incorporate several foundational FL techniques as baselines along with more recent approaches. The methods evaluated are FedAvg (McMahan et al., 2017), FedAdam (Reddi et al., 2021), FedProx (Li et al., 2020), SCAFFOLD (Karimireddy et al., 2020), MOON (Li et al., 2021a), FedPer (Arivazhagan et al., 2019), Ditto (Li et al., 2021b), APFL (Deng et al., 2020), and PerFCL (Zhang et al., 2022). The choice of personalized methods studied in this work is discussed in Appendix A. Two non-FL training setups are also considered. The first, central training, pools all data to train a single model. The second is local training, where each client trains a model solely on local data, resulting in a model for each client with no global information.
3.1 Non-Personalized FL Methods
@@
The original FedAvg algorithm, which applies weighted parameter averaging, remains a strong FL baseline. FedAdam is a recent extension of FedAvg that incorporates server-side first- and second-order momentum estimates and has shown some robustness to data drift. The FedProx and SCAFFOLD methods aim to address issues associated with non-IID datasets by correcting local-weight drift through a regularization term or control variates, respectively. MOON builds on these ideas, applying a contrastive loss to constrain drift in the client models’ feature representations. Each method trains a global set of weights, updated at each server round, shared by all clients.
3.2 Personalized FL Methods
@@
For the personalized FL baselines, each client possesses a unique model. FedPer splits the models into a feature extractor and classifier. The feature extractor is shared among clients and federally trained, while the classifier is unique to each client. Ditto trains a global model with FedAvg and distinct personal models. During client-side training, local model optimization is constrained using a weighted $\ell^{2}$-penalty between the global and local weights. In APFL, predictions are made through a convex combination of twin models. One is federally trained using FedAvg; the other incorporates only local updates. The combination parameter, $\alpha$, is adapted during training using the gradients of the global and local models. Finally, PerFCL extends MOON to consider two feature extractors, one is shared by all clients and the other is specific to each client along with a local classification layer. The feature extractors are constrained using different contrastive losses.
3.3 FENDA-FL
FedPer, APFL, and PerFCL fall under the category of parameter decoupling methods. Such methods reserve a subset of models weights for learning global information and others for local information. In the results below, this type of decoupling provides useful flexibility in clinical settings. For PerFCL, the authors consider a model with parallel global and local feature extractors feeding into a locally trained classifier. The latent spaces of the extractors are constrained with contrastive loss functions that attempt to keep each client’s global latent space close to the aggregated global feature space and to push local features away from these spaces. These losses are weighted with hyper-parameters $\mu$ and $\gamma$. Experiments in the original paper show that PerFCL improves upon MOON and other FL baselines. However, therein, PerFCL is not compared against other personalized methods. Moreover, no ablation studies are considered to quantify the benefit associated with the architecture compared to the loss. As such, we study the effect of removing the contrastive losses.
Removal of the contrastive losses results in an approach that directly extends the FENDA method, originally designed for domain adaptation, to FL. In FENDA, each model associated with a domain has two feature extractors. One is trained on domain-specific data, while the other is shared and trained across all domains. The features are then combined and processed by a domain-specific classification head. The shared feature extractor is free to learn robust, domain-agnostic features, while the domain-specific extractor learns features important in making accurate local predictions.
Assume there are $M$ clients, with local datasets $D_{i}\subset X\times Y$, for $i\in\{1,\dots,M\}$. As with PerFCL, each FENDA-FL client model consists of a global feature extractor, $f_{{\mathbf{w}}}:X\rightarrow\mathbb{R}^{H_{G}}$, a local feature extractor $f_{{\mathbf{w}}_{i}}:X\rightarrow\mathbb{R}^{H_{i}}$, and a local classification head $g_{\theta_{i}}:\mathbb{R}^{(H_{G}+H_{i})}\rightarrow Y$. For a given input sample ${\mathbf{x}}\sim P_{i}(X)$, $f$ and $f_{i}$ map ${\mathbf{x}}$ to representations ${\mathbf{z}}$ and ${\mathbf{z}}_{i}$ which are then concatenated and passed into $g_{\theta_{i}}$ to yield a prediction, $\hat{y}_{i}$, as
$\hat{y}_{i}=g_{\theta_{i}}\left(\begin{bmatrix}{\mathbf{z}}\\ {\mathbf{z}}_{i}\\ \end{bmatrix}\right)=g_{\theta_{i}}\left(\begin{bmatrix}f_{{\mathbf{w}}}({\mathbf{x}})\\ f_{{\mathbf{w}}_{i}}({\mathbf{x}})\\ \end{bmatrix}\right).$
The global extractor weights, ${\mathbf{w}}$, are aggregated across clients via FedAvg at each round. The local extractor and classification head weights, ${\mathbf{w}}_{i}$ and $\theta_{i}$, are exclusively learned through SGD, or its variants, on each client. Appendix I discusses distribution drift robustness and FENDA-FL’s relationship to domain adaptation. As formulated, FENDA-FL is not limited to clinical domains. Nonetheless, in this work, we are primarily interested in the performance of FL approaches for heterogeneous clinical applications.
FENDA-FL has several advantages over PerFCL, beyond a simpler setup. Removal of the contrastive losses eliminates the need to tune two hyper-parameters. Further, it has been observed that similar loss-induced constraints may hinder learning, especially when hyper-parameters are poorly calibrated (Li et al., 2020). Without the contrastive losses, FENDA-FL is also a more flexible approach - global and local modules need not have the same latent space, allowing for the injection of inductive bias through architecture variation. As shown in Appendix F, this is useful in clinical settings to produce optimal performance.
3.4 Federated Checkpointing and Evaluation
Federated evaluation is a nuanced and under-explored process, yet it is a critical component, especially in clinical pipelines (Karargyris et al., 2021; Matsuda et al., 2024). We propose a straightforward checkpointing and evaluation framework for FL models, building on the foundations in (du Terrail et al., 2022a) and better aligning with standard ML model practices. In many works (Li et al., 2020; Karimireddy et al., 2020; Passban et al., 2022), FL training proceeds for a fixed number of server rounds. Thereafter, the final model is evaluated on held-out data, representing either a central or distributed test set. However, as with standard ML training, the model produced at the end of a training trajectory is not always optimal due to phenomena such as over-fitting.
We propose splitting each clients’ dataset into train and validation sets. In the experiments, the split used is 80–20. The validation splits enable two distinct checkpointing approaches. After each round of local training and aggregation, the clients evaluate the model on their local validation sets. In local checkpointing, each client uses their local validation loss to determine whether a new checkpoint should be stored. Alternatively, in global checkpointing, the client losses are aggregated by the server using a weighted average. This average is used to checkpoint a global model shared by all clients. In both cases, the best model is not necessarily the one produced by the final server round. In the local case, clients are free to keep the model they deem best. In the global strategy, all clients share the same model from some point in the server rounds. Personalized FL strategies do not have a single global model and, therefore, only use local checkpointing. When local checkpointing is used, each client loads its respective model to be evaluated on that client’s test data.
Baselines
To evaluate checkpointed models, we advocate for at least three separate baselines to be considered, two of which are part of the FLamby benchmarks. Models trained on central data are a common gold standard for FL performance. In the experiments, a single model is trained on data pooled from each of the clients and checkpointed based on a validation loss. The model is evaluated on each client’s test set and the performance uniformly averaged. In local training, a client trains a model on its own training data. The model is shared with all clients and evaluated on their test data, including that of the original client. This setting mimics the scenario where, for example, a hospital distributes a model to other hospitals. Again, the measured metrics are averaged across clients. An important alternative setting, referred to here as siloed, is one where each client trains its own model and evaluates on its local test data. The results are then averaged across the clients. This simulates a strong baseline where, for example, the problem is important enough that each hospital has collected enough data to train a model. If an FL model does not produce better performance, there is significantly less incentive to participate in collaborative training.
Reproducibility
To quantify stability and smooth randomness in training, five FL training runs with distinct validation splits are performed for each method. Within each run, the performance across clients is averaged. The mean and $95$% confidence intervals for this average are reported. As mentioned above, code to reproduce the experiments is released with this work along with an open-source library.